### v1

@@ 5.2 Matrix representation
$\operatorname{cond}(\operatorname{\mathbf{D}}^{-1}\operatorname{\mathbf{B}})=\left|{\lambda_{\max}(\operatorname{\mathbf{D}}^{-1}\operatorname{\mathbf{B}})}\right|\left|{\lambda_{\max}(\operatorname{\mathbf{B}}^{-1}\operatorname{\mathbf{D}})}\right|\leq\dfrac{\left\lVert{b}\right\rVert\left\lVert{d}\right\rVert}{\widetilde{C}_{B}\widetilde{C}_{D}}.$
@@ 6 Numerical results
This section presents some numerical results to support the theoretical analysis on the Galerkin discretization (16) of the MFIE. We perform numerical experiments for the electromagnetic scattering by two perfectly conducting bodies: a multiply-connected cuboid and a star-based pyramid, see Figure 2. In all experiments, the wave number $\kappa=1$ and the incident electric field is given by
@@
where $\hat{\boldsymbol{x}}$ and $\hat{\boldsymbol{z}}$ stand for the unit vectors along the $x$-axis and $z$-axis, respectively.
In order to examine the convergence of the Galerkin discretization scheme (16), we define the relative error
$\mathrm{err}_{h}:=\dfrac{\left|{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\sim,\operatorname{\mathbf{X}}_{\Gamma}}-\left\lVert{\boldsymbol{u}_{\mathrm{ref}}}\right\rVert_{\sim,\operatorname{\mathbf{X}}_{\Gamma}}}\right|}{\left\lVert{\boldsymbol{u}_{\mathrm{ref}}}\right\rVert_{\sim,\operatorname{\mathbf{X}}_{\Gamma}}},$
where $\boldsymbol{u}_{\mathrm{ref}}$ is a reference solution to (16) computed with a very fine mesh, and the equivalent norm $\left\lVert{\cdot}\right\rVert_{\sim,\operatorname{\mathbf{X}}_{\Gamma}}$ on $\operatorname{\mathbf{X}}_{\Gamma}$ is defined for all $\boldsymbol{u}\in\operatorname{\mathbf{X}}_{\Gamma}$ as follows
$\left\lVert{\boldsymbol{u}}\right\rVert^{2}_{\sim,\operatorname{\mathbf{X}}_{\Gamma}}:=\left\langle{\boldsymbol{u},S_{i\kappa}\overline{\boldsymbol{u}}}\right\rangle_{\times,\Gamma}.$
By means of a triangle inequality, the relative error $\mathrm{err}_{h}$ exhibits the same convergence rate with that stated in Theorem 5.7.
@@ 6.2 Star-based pyramid
where $\omega_{v}$ is the domain on the unit sphere cut out by the tangent cone to $\Gamma$ at a vertex $v$ of $\Gamma$ (see [6, Theorem 8]). On the considered star-based pyramid, the boundary length of the curvilinear polygon $\omega_{v}$ reaches its maximum at the bottom vertex $v_{0}$ of the boundary $\Gamma$, at which 48 edges meet. Particularly, we can find $s^{\ast}\approx 0.228$. Hence, according to Theorem 5.7, only a sub-optimal convergence rate $\mathcal{O}(h^{1.228})$ can be expected on this domain. Figure 3 (right) confirms our convergence analysis. Here, the reference solution $\boldsymbol{u}_{\mathrm{ref}}$ is computed at the meshwidth $h=0.02\mathrm{m}$. Finally, Figure 4 (right) confirms again the well-conditioning property of the proposed Galerkin discretization scheme (16) for the MFIE.
@@ 7 Conclusions
We have provided rigorous numerical analysis of a Galerkin boundary element discretization of the magnetic field integral equation on Lipschitz polyhedra. The continuous variational problem has been shown to be uniquely solvable, provided that the wave number does not belong to the spectrum of the interior Maxwell’s problem. A Galerkin discretization of the variational formulation has been analyzed, which employs Raviart-Thomas boundary elements for the solution space and Buffa-Christiansen elements for the test space. The unique solvability and the well-conditioning of the discrete system have been proven, and the convergence rate of the discretization scheme has been investigated. Some numerical results have also been presented, corroborating the theoretical analysis as well as the effect of the singularities of Maxwell’s solutions on polyhedra to the convergence rate. The analyzed discretization scheme has been widely used in practical applications.

### v2

@@ 5.2 Matrix representation
$\operatorname{cond}(\operatorname{\mathbf{D}}^{-1}\operatorname{\mathbf{B}})=\left|{\lambda_{\max}(\operatorname{\mathbf{D}}^{-1}\operatorname{\mathbf{B}})}\right|\left|{\lambda_{\max}(\operatorname{\mathbf{B}}^{-1}\operatorname{\mathbf{D}})}\right|\leq\dfrac{\left\lVert{b}\right\rVert}{\gamma\alpha}.$
Remark 5.10.
In the paper [12], the authors introduced a family of 3D open-book Lipschitz polyhedra on which the Galerkin discretization of second-kind boundary integral equations for the Laplace and Helmholtz problems do not converge. The negative claims in that paper do not contradict the present results, since in this paper we consider the relevant operators in the natural trace space $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ rather than the space $\operatorname{\mathbf{L}}^{2}(\Gamma)$ as in [12].
Remark 5.11.
The introduction of the purely imaginary wave number $i\kappa^{\prime}$ (and $\kappa^{\prime}_{0}$) is just to check the “quality” of $\Gamma$ whether the proposed Galerkin method for the MFIE converges on that domain. Once Assumption 5.2 is verified, the solvability of the discrete problem (20), the quasi-optimal error estimate and the convergence of numerical solutions are independent of $\kappa^{\prime}_{0}$. However, the bound of the condition number in Theorem 5.8 might be dependent of $\kappa^{\prime}_{0}$ via the constants $\alpha$ and $\gamma$.
6 Verification of Assumption 5.2
Assumption 5.2, along with Assumption 1, gives a sufficient condition for the unique solvability of the discrete problem (20). However, verifying Assumption 5.2 for each domain is not straightforward. In this section, we propose a numerical scheme to approximately compute the constants $\alpha$ and $\beta$, then verify this assumption in practice.
We first evaluate the two terms in the denominator of (19). For $\kappa^{\prime}>0$, let $\operatorname{\mathbf{S}}_{\kappa^{\prime}}$ be the Galerkin matrix of the $\operatorname{\mathbf{X}}_{\kappa^{\prime}}$-inner product $\left({\cdot,\cdot}\right)_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}:\operatorname{\mathbf{U}}_{h}\times\operatorname{\mathbf{U}}_{h}\to\mathbb{C}$. Then, the $\operatorname{\mathbf{X}}_{\kappa^{\prime}}$-norm of $\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}$ can be expressed as the $\textbf{{l}}^{2}$-norm of its coefficient vector $\widehat{\boldsymbol{u}}_{h}$ as follows
(28) $\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}=\sqrt{\widehat{\boldsymbol{u}}_{h}^{\mathsf{T}}\operatorname{\mathbf{S}}_{\kappa^{\prime}}\widehat{\boldsymbol{u}}_{h}}=\left\lVert{\operatorname{\mathbf{S}}_{\kappa^{\prime}}^{1/2}\widehat{\boldsymbol{u}}_{h}}\right\rVert_{\textbf{{l}}^{2}}.$
In order to evaluate $\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{v}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$, with $\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}$, we approximate $S^{-1}_{i\kappa^{\prime}}\boldsymbol{v}_{h}$ by its projection $\widetilde{\boldsymbol{w}}_{h}\in\widetilde{\operatorname{\mathbf{U}}}_{h}$ with respect to the $\operatorname{\mathbf{X}}_{\kappa^{\prime}}$-inner product, i.e., $\widetilde{\boldsymbol{w}}_{h}\in\widetilde{\operatorname{\mathbf{U}}}_{h}$ is the unique solution to the following equation
(29) $\left\langle{\widetilde{\boldsymbol{w}}_{h},S_{i\kappa^{\prime}}\overline{\widetilde{\boldsymbol{\varphi}}_{h}}}\right\rangle_{\times,\Gamma}=-\left\langle{\boldsymbol{v}_{h},\overline{\widetilde{\boldsymbol{\varphi}}_{h}}}\right\rangle_{\times,\Gamma},\qquad\qquad\quad\forall\widetilde{\boldsymbol{\varphi}}_{h}\in\widetilde{\operatorname{\mathbf{U}}}_{h}.$
Here, the subspace $\widetilde{\operatorname{\mathbf{U}}}_{h}$ consists of all Raviart-Thomas boundary elements defined on the barycentric refinement $\widetilde{\Gamma}_{h}$ of $\Gamma_{h}$. Please note that $\operatorname{\mathbf{U}}_{h}\subset\widetilde{\operatorname{\mathbf{U}}}_{h},\operatorname{\mathbf{V}}_{h}\subset\widetilde{\operatorname{\mathbf{U}}}_{h}$, and $\dim\widetilde{\operatorname{\mathbf{U}}}_{h}=6\dim\operatorname{\mathbf{U}}_{h}=6\dim\operatorname{\mathbf{V}}_{h}=6N$. Let $\widetilde{\operatorname{\mathbf{S}}}_{\kappa^{\prime}}$ and $\operatorname{\mathbf{G}}$ be the Galerkin matrices of the sesquilinear forms defined by the left-hand side and right-hand side of (29), respectively. Then, the coefficient vector $\widehat{\boldsymbol{w}}_{h}$ of $\widetilde{\boldsymbol{w}}_{h}$ with respect to the basis of $\widetilde{\operatorname{\mathbf{U}}}_{h}$ is determined by
$\widehat{\boldsymbol{w}}_{h}=\widetilde{\operatorname{\mathbf{S}}}_{\kappa^{\prime}}^{-1}\operatorname{\mathbf{G}}\widehat{\boldsymbol{v}}_{h},$
where $\widehat{\boldsymbol{v}}_{h}$ is the coefficient vector of $\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}$ with respect to the basis $\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{N}\}$. Thus, $\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{v}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}$ is approximated by
(30) $\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{v}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\approx\left\lVert{\widetilde{\boldsymbol{w}}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}=\sqrt{\widehat{\boldsymbol{w}}_{h}^{\mathsf{T}}\widetilde{\operatorname{\mathbf{S}}}_{\kappa^{\prime}}\widehat{\boldsymbol{w}}_{h}}=\left\lVert{\left({\operatorname{\mathbf{G}}^{\mathsf{T}}\widetilde{\operatorname{\mathbf{S}}}_{\kappa^{\prime}}^{-1}\operatorname{\mathbf{G}}}\right)^{1/2}\widehat{\boldsymbol{v}}_{h}}\right\rVert_{\textbf{{l}}^{2}}.$
Let $\operatorname{\mathbf{M}}_{\kappa^{\prime}}:=\operatorname{\mathbf{G}}^{\mathsf{T}}\widetilde{\operatorname{\mathbf{S}}}_{\kappa^{\prime}}^{-1}\operatorname{\mathbf{G}}$. The matrix $\operatorname{\mathbf{M}}_{\kappa^{\prime}}$ is positive-definite, as $\widetilde{\operatorname{\mathbf{S}}}_{\kappa^{\prime}}$ is positive-definite and $\operatorname{\mathbf{G}}$ is injective. Based on the matrix representations (28) and (30), the left-hand side of the inf-sup condition (19) can be expressed in the following form
$\inf_{\boldsymbol{u}_{h}\in\operatorname{\mathbf{U}}_{h}}\sup_{\boldsymbol{v}_{h}\in\operatorname{\mathbf{V}}_{h}}\dfrac{\left|{\left\langle{\boldsymbol{u}_{h},\overline{\boldsymbol{v}_{h}}}\right\rangle_{\times,\Gamma}}\right|}{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}\left\lVert{S^{-1}_{i\kappa^{\prime}}\boldsymbol{v}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa^{\prime}}}}$ $\approx\inf_{\widehat{\boldsymbol{u}}_{h}\in\mathbb{C}^{N}}\sup_{\widehat{\boldsymbol{v}}_{h}\in\mathbb{C}^{N}}\dfrac{\left|{\widehat{\boldsymbol{v}}_{h}^{\mathsf{T}}\operatorname{\mathbf{D}}\widehat{\boldsymbol{u}}_{h}}\right|}{\left\lVert{\operatorname{\mathbf{S}}_{\kappa^{\prime}}^{1/2}\widehat{\boldsymbol{u}}_{h}}\right\rVert_{\textbf{{l}}^{2}}\left\lVert{\operatorname{\mathbf{M}}_{\kappa^{\prime}}^{1/2}\widehat{\boldsymbol{v}}_{h}}\right\rVert_{\textbf{{l}}^{2}}}$
$=\inf_{\widehat{\boldsymbol{\xi}}_{h}\in\mathbb{C}^{N}}\sup_{\widehat{\boldsymbol{\psi}}_{h}\in\mathbb{C}^{N}}\dfrac{\left|{\widehat{\boldsymbol{\psi}}_{h}^{\mathsf{T}}\operatorname{\mathbf{M}}_{\kappa^{\prime}}^{-\mathsf{T}/2}\operatorname{\mathbf{D}}\operatorname{\mathbf{S}}_{\kappa^{\prime}}^{-1/2}\widehat{\boldsymbol{\xi}}_{h}}\right|}{\left\lVert{\widehat{\boldsymbol{\xi}}_{h}}\right\rVert_{\textbf{{l}}^{2}}\left\lVert{\widehat{\boldsymbol{\psi}}_{h}}\right\rVert_{\textbf{{l}}^{2}}},$
where $\operatorname{\mathbf{M}}_{\kappa^{\prime}}^{-\mathsf{T}/2}=\left({\operatorname{\mathbf{M}}_{\kappa^{\prime}}^{-1/2}}\right)^{\mathsf{T}}$. Hence, for a sufficiently small $h>0$, the constant $\alpha$ in (19) can be approximated by
$\alpha\approx\left|{\lambda_{\min}\left({\operatorname{\mathbf{M}}_{\kappa^{\prime}}^{-\mathsf{T}/2}\operatorname{\mathbf{D}}\operatorname{\mathbf{S}}_{\kappa^{\prime}}^{-1/2}}\right)}\right|.$
Please bear in mind that the space $\mathbb{C}^{N}$ is now equipped with the Euclidean $\textbf{{l}}^{2}$-norm. Analogously, taking into account the property (9), $\beta$ can be estimated as
(31) $\beta\approx\left|{\lambda_{\max}\left({\operatorname{\mathbf{M}}_{\kappa^{\prime}}^{-\mathsf{T}/2}\operatorname{\mathbf{K}}_{\kappa^{\prime}}\operatorname{\mathbf{S}}_{\kappa^{\prime}}^{-1/2}}\right)}\right|,$
where $\operatorname{\mathbf{K}}_{\kappa^{\prime}}$ is the Galerkin matrix associated with the operator $C_{i\kappa^{\prime}}$, i.e.,
${[\operatorname{\mathbf{K}}_{\kappa^{\prime}}]}_{mn}:=\left\langle{\boldsymbol{u}_{n},C_{i\kappa^{\prime}}\overline{\boldsymbol{v}_{m}}}\right\rangle_{\times,\Gamma},\qquad\quad m,n=1,2,\ldots,N.$
Finally, we can verify Assumption 5.2 for the boundary $\Gamma$ by checking if the condition $\tfrac{\alpha}{2}-\beta>0$ is fulfilled by some $\kappa^{\prime}>0$.
Remark 6.1.
Another scheme to approximate the constant $\beta$ is to replace the Galerkin matrix $\operatorname{\mathbf{K}}_{\kappa^{\prime}}$ by its counterpart $\widetilde{\operatorname{\mathbf{K}}}_{\kappa^{\prime}}$ tested with $\operatorname{\mathbf{U}}_{h}$, i.e.,
${[\widetilde{\operatorname{\mathbf{K}}}_{\kappa^{\prime}}]}_{mn}:=\left\langle{\boldsymbol{u}_{n},C_{i\kappa^{\prime}}\overline{\boldsymbol{u}_{m}}}\right\rangle_{\times,\Gamma},\qquad\qquad m,n=1,2,\ldots,N.$
Then, the matrices $\operatorname{\mathbf{G}}$ and $\operatorname{\mathbf{M}}_{\kappa^{\prime}}$ must also be changed accordingly (the latter is denoted by $\widetilde{\operatorname{\mathbf{M}}}_{\kappa^{\prime}}$). The advantage of this approach is the symmetry of the matrix $\widetilde{\operatorname{\mathbf{K}}}_{\kappa^{\prime}}$, due to (8). As a consequence, all eigenvalues of $\widetilde{\operatorname{\mathbf{M}}}_{\kappa^{\prime}}^{-\mathsf{T}/2}\widetilde{\operatorname{\mathbf{K}}}_{\kappa^{\prime}}\operatorname{\mathbf{S}}_{\kappa^{\prime}}^{-1/2}$ are real, and its maximal eigenvalue $\left|{\lambda_{\max}}\right|$ might give a better approximation of $\beta$. The disadvantage is that we have to additionally compute the matrix $\widetilde{\operatorname{\mathbf{M}}}_{\kappa^{\prime}}$ and its square root inverse, which is quite expensive. From authors’ experience, these two approaches produce similar results. Therefore, the scheme (31) is recommended.
Remark 6.2.
From authors’ experience, Assumption 5.2 is satisfied for typical geometries in practical applications, usually with larger $\kappa^{\prime}>0$ for less regular domains (see Figure 3 for an example). When discretizing the integral operators with larger $\kappa^{\prime}$, finer meshes and/or more quadrature points are required to accurately compute the non-zero entries of the corresponding Galerkin matrices. As a result, the verification step may cost more than the computational cost of solving the MFIE itself. Therefore, verifying Assumption 5.2 beforehand is usually not desirable in practice.
@@ 7 Numerical results
This section presents some numerical results to support the analysis of the Galerkin discretization (20) of the MFIE. We perform numerical experiments for the electromagnetic scattering by two perfectly conducting bodies: a multiply-connected cuboid and a star-based pyramid, see Figure 2. The verification of Assumption 5.2 for these domains are firstly performed. The approximated values of $\alpha$ and $\beta$ with respect to different $\kappa^{\prime}$ are depicted in Figure 3, which clearly shows that there are some values of $\kappa^{\prime}$ fulfilling Assumption 5.2 for each domain (the values at which the blue curve is above the red one).
Figure 3: Approximated constants $\alpha$ and $\beta$ for the verification of Assumption 5.2. Left: A cuboid of dimension $\mathrm{1\mathrm{m}\times 1\mathrm{m}\times 0.25\mathrm{m}}$ with a concentric cuboid hole of dimension $\mathrm{0.5\mathrm{m}\times 0.5\mathrm{m}\times 0.25\mathrm{m}}$. Right: A pyramid of height $0.5\mathrm{m}$ and a 24-pointed star base, whose vertices lie on two circles of radius $1\mathrm{m}$ and $0.3\mathrm{m}$. For each domain, there are some $\kappa^{\prime}$ at which Assumption 5.2 is satisfied.
In all further experiments, the wave number $\kappa=1$ and the incident electric field is given by
@@
where $\hat{\boldsymbol{x}}$ and $\hat{\boldsymbol{z}}$ stand for the unit vectors along the $x$-axis and $z$-axis, respectively. In order to examine the convergence of the Galerkin discretization scheme (20), we define the relative error
$\mathrm{err}_{h}:=\dfrac{\left|{\left\lVert{\boldsymbol{u}_{h}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa}}-\left\lVert{\boldsymbol{u}_{\mathrm{ref}}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa}}}\right|}{\left\lVert{\boldsymbol{u}_{\mathrm{ref}}}\right\rVert_{\operatorname{\mathbf{X}}_{\kappa}}},$
where $\boldsymbol{u}_{\mathrm{ref}}$ is a reference solution to (20) computed with a very fine mesh. By means of a triangle inequality, the relative error $\mathrm{err}_{h}$ exhibits the same convergence rate with that stated in Theorem 5.6.
@@ 7.2 Star-based pyramid
where $\omega_{v}$ is the domain on the unit sphere cut out by the tangent cone to $\Gamma$ at a vertex $v$ of $\Gamma$ (see [7, Theorem 8]). On the considered star-based pyramid, the boundary length of the curvilinear polygon $\omega_{v}$ reaches its maximum at the bottom vertex $v_{0}$ of the boundary $\Gamma$, at which 48 edges meet. Particularly, we can find $s^{\ast}\approx 0.228$. Hence, according to Theorem 5.6, only a sub-optimal convergence rate $\mathcal{O}(h^{1.228})$ can be expected on this domain. Figure 4 (right) confirms our convergence analysis. Here, the reference solution $\boldsymbol{u}_{\mathrm{ref}}$ is computed at the meshwidth $h=0.02\mathrm{m}$. Finally, Figure 5 (right) confirms again the well-conditioning property of the proposed Galerkin discretization scheme (20) for the MFIE.
@@ 8 Conclusions
We have provided rigorous analysis of a Galerkin boundary element discretization of the magnetic field integral equation on Lipschitz polyhedra. The continuous variational problem has been shown to be uniquely solvable, provided that the wave number does not belong to the spectrum of the interior Maxwell’s problem. A Petrov-Galerkin discretization of the variational formulation employing Raviart-Thomas boundary elements for the solution space and Buffa-Christiansen elements for the test space has been analyzed. Under a stability condition, the solvability and the well-conditioning of the discrete system have been proven, and the convergence rate of the discretization scheme has been investigated. A numerical scheme to verify the stability condition has been proposed, showing that this condition is unproblematic in practice. Some numerical results have also been presented, corroborating the theoretical analysis as well as the effect of the singularities of Maxwell’s solutions on polyhedra to the convergence rate. The analyzed discretization scheme has been widely used in practical applications.